high pressure steam enters a turbine with a velocity of 38.0 meters/sec and leaves the exhaust trunk with a velocity of 305 meters/sec. calculate the kinetic energy of the steam in KJ/kg of the steam in KJ/kg and N-m/kg at a. at the entrance to the turbine and b. the exhaust trunk
1
Expert's answer
2020-03-02T09:12:47-0500

As per the given question,

Velocity of the steam at the entry point "(v_1)" = 38 m/sec

Velocity of the steam at the exhaust trunk "(v_2)" =305 m/sec

We know that density of the steam is =0.6 "kg\/m^3"

a)

So, kinetic energy of the steam at the entrance to the turbine="\dfrac{m v^2}{2}"

if m=1 kg and energy per kg will be "=\dfrac{mv^2}{m}=\dfrac{v^2}{2}"

="\dfrac{38^2}{2}=722J/kg"

or "0.722 K J\/kg"

b)

Kinetic energy at the exhaust trunk,

"=\dfrac{m v_2^2}{2}"

if m=1 kg and energy per kg will be "=\dfrac{mv_2^2}{m}=\dfrac{v_2^2}{2}"

"=\dfrac{305^2}{2}=" 46512.5 J/kg

or 46.5125 KJ/kg